OpenStudy (anonymous):

Find the point of intersection of the following system of equations. y = x + 9 y = -2x + 6

OpenStudy (igreen):

It's asking for the solution. We can plug in x + 9 for 'y' in the 2nd equation to solve for 'x'. y = -2x + 6 x + 9 = -2x + 6 Subtract 9 to both sides, what's 6 - 9?

OpenStudy (anonymous):

-3

OpenStudy (igreen):

Yes, so we have: x = -2x - 3 Add 2x to both sides, what's x + 2x?

OpenStudy (anonymous):

3x?

OpenStudy (igreen):

Yep, so we have: 3x = -3 Now divide 3 to both sides, what's -3 / 3?

OpenStudy (anonymous):

-1

OpenStudy (igreen):

Yes, so the x-value of our intersection is -1, we can plug this into any of the two equations to get our y-value. y = x + 9 y = -1 + 9 Add -1 + 9

OpenStudy (anonymous):

8

OpenStudy (igreen):

Yes, so the y-value of our intersection is 8. So our full point is (-1, 8).

OpenStudy (anonymous):

thank you